# Category $\mathcal{O}$ for hybrid quantum groups and non-commutative Springer resolutions

### v3

@@ 5.2. Truncated categories
(5.5) $D^{b}\mathrm{Coh}^{G}(\widehat{\mathrm{St}})\simeq\widehat{D}_{\check{I}_{0}}(\widetilde{\mathcal{F}{l}}),$
where $\widehat{D}_{\check{I}_{0}}(\widetilde{\mathcal{F}{l}})$ is a completion of the category of $\check{I}_{0}$-equivariant monodromic complexes on $\widetilde{\mathcal{F}{l}}$. There are standard and costandard sheaves $j_{x!}$, $j_{x*}$ in $\widehat{D}_{\check{I}_{0}}(\widetilde{\mathcal{F}{l}})$ associated to the embedding $j_{x}$ from the Schubert cell labelled by $x\in W_{\mathrm{ex}}$ to $\widetilde{\mathcal{F}{l}}$. The equivalence (5.5) sends ${\mathcal{O}}_{\tilde{{\mathfrak{g}}}_{\Delta}}$ to $j_{e!}$, and in general ${\mathcal{O}}_{\tilde{{\mathfrak{g}}}_{\Delta}}(w_{0}\lambda)$ to the Wakimoto sheaf ${\mathcal{J}}_{\lambda}$ (see [8, §3.3]), for any $\lambda\in\Lambda$ (In loc. cit. the line bundle labelled by $\lambda$ is sent to ${\mathcal{J}}_{\lambda}$, and here our correspondence is due to the different convention that ${\mathcal{O}}_{{\mathcal{B}}}(\lambda)$ is semi-ample if and only if $\lambda$ is anti-dominant).
We view $\widehat{\mathrm{St}}$ as the base change of $\pi:\tilde{{\mathfrak{g}}}\rightarrow{\mathfrak{g}}$ (as the second factor) by $\tilde{{\mathfrak{g}}}\times_{{\mathfrak{t}}/W}({\mathfrak{t}}/W)_{\hat{0}}\rightarrow{\mathfrak{g}}$ (as the first factor), then by Theorem 3.1 there are functors ${\mathbb{T}}^{c}_{s}$, $s\in{\mathrm{I}}$ on $D^{b}\mathrm{Coh}^{G}(\widehat{\mathrm{St}})$. By [8] the equivalence (5.5) intertwines the functor ${\mathbb{T}}^{c}_{s}$ with the convolution with $j_{s!}$ from the right. Therefore, by (4.15) the composed equivalence
$D^{b}({\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B})\simeq\widehat{D}_{\check{I}_{0}}(\widetilde{\mathcal{F}{l}})$
@@
Let $\mu$ be dominant enough, then we have ${\mathcal{J}}_{\lambda+\mu}=j_{t(\lambda+\mu)*}$ (see [8, §3.3]). Since $t(\lambda)$ is maximal in the Bruhat order on $t(\lambda)W$ if $\lambda$ is dominant, we have ${\mathcal{J}}_{\lambda+\mu}*j_{w!}=j_{t(\lambda+\mu)*}*j_{w!}=j_{t(\lambda+\mu)w*}$. Note that
@@
which is nonzero only if $t(\lambda^{\prime}+\mu)w^{\prime}\leq t(\lambda+\mu)w$ in the Bruhat order. Since $t(\lambda^{\prime}+\mu)w^{\prime}$ and $t(\lambda+\mu)w$ translate the fundamental alcove to the dominant alcoves, the Bruhat order on them coincides with the semi-infinite order (see e.g. [39, Claim 4.14]). We deduce that
@@ 5.3. Proof of the main result
Recall the induction functor
@@
There are natural isomorphisms
${\mathbb{V}}_{r}\circ\mathrm{ind}_{r}\simeq\mathrm{ind}_{r}\circ{\mathbb{V}}^{b}_{r}\ ,\quad\text{and}\quad{\mathbb{V}}_{c}\circ\mathrm{ind}_{c}\simeq\mathrm{ind}_{c}\circ{\mathbb{V}}^{b}_{c}.$
@@
where the second isomorphism is by tensor formula. Since the action of $Z_{\mathrm{HC}}$ on $U^{\mathrm{hb}}_{\zeta}\otimes_{{\mathfrak{U}}^{b}_{\zeta}}M$ coincides with the one induced by the $Z_{\mathrm{HC}}$-action on $M$, the equality ${\mathbb{V}}_{r}\circ\mathrm{ind}_{r}(M)=\mathrm{ind}_{r}\circ{\mathbb{V}}^{b}_{r}(M)$ is also an isomorphism of $S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]$-modules. It implies the first natural isomorphism. The second isomorphism is straight forward. ∎
@@
We firstly prove the assertion for ${\mathbb{V}}_{c}$ in a special case: let $M\in{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$ admitting Verma flags, and let $Q=\mathrm{ind}_{c}(M^{\prime})^{\leq\nu}$ with $M^{\prime}\in{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ admitting Verma flags. We show that the natural map
@@
where the horizontal isomorphisms follow from the adjunctions for $(-)^{\leq\nu}$ and $\mathrm{ind}_{c}$, and the equalities
${\mathbb{V}}_{c}(\mathrm{ind}_{c}(M^{\prime})^{\leq\nu})={\mathbb{V}}_{c}(\mathrm{ind}_{c}(M^{\prime}))^{\leq\nu}\quad\text{and}\quad{\mathbb{V}}_{c}(\mathrm{ind}_{c}(M^{\prime}))=\mathrm{ind}_{c}({\mathbb{V}}^{b}_{c}(M^{\prime}))$
by Lemma 5.8 and Lemma 5.10. It is now equivalent to show that the right vertical map is an isomorphism, which is true by Proposition 4.13.
In general, let $M_{1},M_{2}\in{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$ admitting Verma flags. By Lemma 5.7, there is a resolution $Q_{1}\rightarrow Q_{2}\rightarrow M_{1}\rightarrow 0$ with $Q_{i}=\mathrm{ind}_{c}(P_{i})^{\leq\nu}$ for some projective objects $P_{i}$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$. We have a commutative diagram with exact rows
00{0}${\mathrm{Hom}(M_{1},M_{2})}$${\mathrm{Hom}(Q_{2},M_{2})}$${\mathrm{Hom}(Q_{1},M_{2})}$00{0}${\mathrm{Hom}({\mathbb{V}}_{c}M_{1},{\mathbb{V}}_{c}M_{2})}$${\mathrm{Hom}({\mathbb{V}}_{c}Q_{2},{\mathbb{V}}_{c}M_{2})}$${\mathrm{Hom}({\mathbb{V}}_{c}Q_{1},{\mathbb{V}}_{c}M_{2}),}$
@@
For any commutative Noetherian $S$-algebra $R$, the equivalence $\mathfrak{G}$ induces an equivalence of the categories on both sides after base changed to $R$. In particular, we have
@@
${\mathfrak{F}}:\ \mathrm{Coh}^{G}(\pi^{*}{\mathrm{A}}_{({\mathfrak{t}}/W)_{\hat{0}}}){\ \xrightarrow{\sim}\ }{\mathcal{O}}_{S}^{0},$
@@
${\mathfrak{F}}\big{(}{\mathbb{T}}^{c}_{w}({\mathcal{E}}^{\vee}_{\hat{0}})\otimes{\mathcal{O}}_{\tilde{{\mathfrak{g}}}}(\nu)\big{)}=M(w_{0}w^{-1}\bullet 0+l\nu)_{S},\quad\forall w\in W,\ \forall\nu\in\Lambda.$
@@
Let $P$ be a projective module in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$, and set ${\mathcal{P}}=({\mathfrak{F}}^{b})^{-1}(P)$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$. Set $Q=\mathrm{ind}_{r}(P)$ and ${\mathcal{Q}}=\mathrm{ind}_{c}({\mathcal{P}})$. We have the following isomorphisms by Lemma 5.9 and Lemma 5.10,
${\mathbb{V}}_{r}(Q^{\leq l\nu})=({\mathbb{V}}_{r}Q)^{\leq l\nu}=({\mathbb{V}}_{r}\circ\mathrm{ind}_{r}P)^{\leq l\nu}=(\mathrm{ind}_{r}\circ{\mathbb{V}}^{b}_{r}P)^{\leq l\nu},$
and similarly ${\mathbb{V}}_{c}({\mathcal{Q}}^{\leq\nu})=(\mathrm{ind}_{c}\circ{\mathbb{V}}_{c}^{b}{\mathcal{P}})^{\leq\nu}$. By Proposition 4.6 and Corollary 4.7, we have
$\mathfrak{G}(\mathrm{ind}_{c}\circ{\mathbb{V}}_{c}^{b}{\mathcal{P}})^{\leq\nu}=(\mathfrak{G}\circ\mathrm{ind}_{c}\circ{\mathbb{V}}_{c}^{b}{\mathcal{P}})^{\leq l\nu}=(\mathrm{ind}_{r}\circ\mathfrak{G}^{b}\circ{\mathbb{V}}_{c}^{b}{\mathcal{P}})^{\leq l\nu}=(\mathrm{ind}_{r}\circ{\mathbb{V}}_{r}^{b}P)^{\leq l\nu}.$
It follows that $\mathfrak{G}\big{(}{\mathbb{V}}_{c}({\mathcal{Q}}^{\leq\nu})\big{)}={\mathbb{V}}_{r}(Q^{\leq l\nu})$. Hence we obtain identifications
${\mathfrak{G}}\circ{\mathbb{V}}_{c}\big{(}\mathrm{Proj}({\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu})\big{)}={\mathbb{V}}_{r}\big{(}\mathrm{Proj}({\mathcal{O}}^{0,\leq l\nu}_{S})\big{)},\quad\nu\in\Lambda$
@@
$\mathrm{Proj}({\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}){\ \xrightarrow{\sim}\ }\mathrm{Proj}({\mathcal{O}}^{0,\leq l\nu}_{S}),\quad\forall\nu\in\Lambda,$
@@
${\mathfrak{F}}:\ {\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}{\ \xrightarrow{\sim}\ }{\mathcal{O}}^{0}_{S}.$
Then we identify ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}=\mathrm{Coh}^{G}(\pi^{*}{\mathrm{A}}_{({\mathfrak{t}}/W)_{\hat{0}}})$ by induction from $B$ to $G$.
@@
[…]
Hence ${\mathfrak{F}}({\mathcal{M}}_{x})$ ($x\in W_{\mathrm{ex}}$) is isomorphic to $M(x\bullet_{l}0)_{S}$ as a $\Lambda$-graded ${\mathfrak{U}}^{b}_{\zeta}\otimes S$-module. It forces that ${\mathfrak{F}}({\mathcal{M}}_{x})\cong M(x\bullet_{l}0)_{S}$ as $\Lambda$-graded $U^{\mathrm{hb}}_{\zeta}\otimes S$-modules. Now the second assertion follows from (4.15). ∎
@@
The assumption that $l$ is a power of prime is used in the results in §3.2 concerning the localization theorem for ${\mathfrak{U}}_{\zeta}$. However, it is expected that this assumption is not necessary (see a proof for type $A$ in [41]). For Theorem 5.13, one may remove this assumption by showing that the equivalence class of the category ${\mathcal{O}}^{0}_{S}$ is independent of $\zeta$ (note that similar independence for small quantum groups appeared in [1].)
Let $\pi_{{}_{\widetilde{{\mathcal{N}}}}}:\widetilde{{\mathcal{N}}}=G\times^{B}{\mathfrak{n}}\rightarrow{\mathfrak{g}}$ be the Springer resolution. We set ${\mathrm{A}}^{\prime}$ be the specialization of ${\mathrm{A}}$ at $0\in{\mathfrak{t}}$, which satisfies an equivalence [12]
$D^{b}\mathrm{Coh}(\widetilde{{\mathcal{N}}})\ \simeq\ D^{b}({\mathrm{A}}^{\prime}\text{-}\mathrm{mod}),$
and therefore is called the non-commutative Springer resolution. Recall that ${\mathrm{A}}_{0}$ is the specialization of ${\mathrm{A}}$ at $0\in{\mathfrak{g}}$. We abbreviate ${\mathcal{O}}^{\chi_{0}}_{\mathbb{C}}$ as the full subcategory of ${\mathcal{O}}_{\mathbb{C}}$ of modules vanished by $\ker\chi_{0}\subset Z_{\mathrm{HC}}$.
@@
(5.6) $\mathrm{Coh}^{G}(\pi_{{}_{\widetilde{{\mathcal{N}}}}}^{*}{\mathrm{A}})={\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}{\ \xrightarrow{\sim}\ }{\mathcal{O}}_{\mathbb{C}}^{0},$
@@
$\mathrm{Coh}^{G}(\pi_{{}_{\widetilde{{\mathcal{N}}}}}^{*}{\mathrm{A}}^{\prime}){\ \xrightarrow{\sim}\ }{\mathcal{O}}_{\mathbb{C}}^{{\chi_{0}}},\quad\text{and}\quad{\mathrm{A}}_{0}\text{-}\mathrm{mod}^{B}{\ \xrightarrow{\sim}\ }{\mathcal{O}}^{u_{\zeta}U^{+}_{\zeta},0}_{\mathbb{C}}.$

### v4

@@ 5.2.3. Truncation by $\leq^{\frac{\infty}{2}}$
[…]
@@
Let $\mu$ be dominant enough, then we have ${\mathcal{J}}_{\lambda+\mu}=j_{t(\lambda+\mu)*}$ (see [14, §3.3]). Since $t(\lambda)$ is maximal in the Bruhat order on $t(\lambda)W$ if $\lambda$ is dominant, we have ${\mathcal{J}}_{\lambda+\mu}*j_{w!}=j_{t(\lambda+\mu)*}*j_{w!}=j_{t(\lambda+\mu)w*}$. Note that
@@
which is nonzero only if $t(\lambda^{\prime}+\mu)w^{\prime}\leq t(\lambda+\mu)w$ in the Bruhat order. Since $t(\lambda^{\prime}+\mu)w^{\prime}$ and $t(\lambda+\mu)w$ translate the fundamental alcove to the dominant alcoves, the Bruhat order on them coincides with the semi-infinite order (see e.g. [37, Claim 4.14]). We deduce that
@@ 5.3. Proof of the main result
In this section, we prove our main result.
@@
There is a natural isomorphism
(5.10) ${\mathbb{V}}_{r}\circ\mathrm{ind}_{r}\simeq\mathrm{ind}_{r}\circ{\mathbb{V}}^{b}_{r}$
of functor from ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$ to $U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,-\rho}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}$, and a natural isomorphism
(5.11) ${\mathbb{V}}_{c}\circ\mathrm{ind}_{c}\simeq\mathrm{ind}_{c}\circ{\mathbb{V}}^{b}_{c}$
of functors from ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ to $({\mathbb{C}}[{\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]\rtimes U{\mathfrak{n}})\text{-}\mathrm{mod}^{T}$.
@@
where the second isomorphism is by tensor formula. Since the action of $Z_{\mathrm{HC}}$ on $U^{\mathrm{hb}}_{\zeta}\otimes_{{\mathfrak{U}}^{b}_{\zeta}}M$ coincides with the one induced by the $Z_{\mathrm{HC}}$-action on $M$, the equality ${\mathbb{V}}_{r}\circ\mathrm{ind}_{r}(M)=\mathrm{ind}_{r}\circ{\mathbb{V}}^{b}_{r}(M)$ is also an isomorphism of $S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]$-modules. It implies (5.10).
The isomorphism (5.11) is straight forward to check. ∎
@@
We firstly prove the assertion for ${\mathbb{V}}_{c}$ in a special case: let $M\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$ admitting Verma flags, and let $Q=\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(\mathrm{ind}_{c}(M^{\prime}))$ with $M^{\prime}\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ admitting Verma flags. We show that the natural map
@@
$\mathrm{Hom}(Q,M)$$\mathrm{Hom}(\mathrm{ind}_{c}(M^{\prime}),M)$$\mathrm{Hom}(M^{\prime},M)$$\mathrm{Hom}({\mathbb{V}}_{c}Q,{\mathbb{V}}_{c}M)$$\mathrm{Hom}({\mathbb{V}}_{c}(\mathrm{ind}_{c}(M^{\prime})),{\mathbb{V}}_{c}M)$$\mathrm{Hom}({\mathbb{V}}_{c}^{b}M^{\prime},{\mathbb{V}}_{c}^{b}M),$$\scriptstyle{{\mathbb{V}}_{c}}\scriptstyle{\simeq}$$\scriptstyle{{\mathbb{V}}_{c}}\scriptstyle{\simeq}$$\scriptstyle{{\mathbb{V}}_{c}^{b}}\scriptstyle{\simeq}\scriptstyle{\simeq}$
where the horizontal isomorphisms follow from the adjunctions for $\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}$, $\tau_{c}^{\leq\nu}$ and $\mathrm{ind}_{c}$, and the natural isomorphisms (5.6) and (5.11). It is now equivalent to show that the right vertical map is an isomorphism, which is true by Proposition 4.13.
In general, let $M_{1},M_{2}\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$ admitting Verma flags. By Lemma 5.18, there is a resolution $Q_{2}\rightarrow Q_{1}\rightarrow M_{1}\rightarrow 0$ with $Q_{i}=\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(\mathrm{ind}_{c}(P_{i}))$ for some projective objects $P_{i}$ in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$. We have a commutative diagram with exact rows
@@
By (2.5), for any commutative Noetherian $S$-algebra $R$, the equivalence $\mathfrak{G}$ induces an equivalence of the categories on both sides after base changed to $R$. In particular, we have
@@
${\mathfrak{F}}:\ \mathrm{Coh}^{G}(\pi^{*}{\mathsf{A}}_{({\mathfrak{t}}/W)_{\hat{0}}}){\ \xrightarrow{\sim}\ }{\mathcal{O}}_{S}^{0},$
@@
${\mathfrak{F}}\big({\mathbb{T}}^{c}_{w}({\mathcal{E}}^{\vee}_{\hat{0}})\otimes{\mathcal{O}}_{\widetilde{{\mathfrak{g}}}}(\nu)\big)=M(w_{0}w^{-1}\bullet 0+l\nu)_{S},\quad\forall w\in W,\ \forall\nu\in\Lambda.$
@@
Let $P$ be a projective module in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$, and set ${\mathcal{P}}=({\mathfrak{F}}^{b})^{-1}(P)$ in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$. Set $Q=\mathrm{ind}_{r}(P)$ and ${\mathcal{Q}}=\mathrm{ind}_{c}({\mathcal{P}})$. We have the following isomorphisms by (5.2) and (5.10),
${\mathbb{V}}_{r}(\tau^{\leq^{\frac{\infty}{2}}t(\nu)}_{r}Q)=\tau^{\leq-\rho+l\nu}_{r}({\mathbb{V}}_{r}Q)=\tau^{\leq-\rho+l\nu}_{r}(\mathrm{ind}_{r}{\mathbb{V}}^{b}_{r}(P)),$
and similarly by (5.6) and (5.11) we have
${\mathbb{V}}_{c}(\tau^{\leq^{\frac{\infty}{2}}t(\nu)}_{c}{\mathcal{Q}})=\tau^{\leq\nu}_{c}({\mathbb{V}}_{c}{\mathcal{Q}})=\tau^{\leq\nu}_{c}(\mathrm{ind}_{c}{\mathbb{V}}^{b}_{c}({\mathcal{P}})).$
By (4.11), (4.10) and (4.15), we have isomorphisms
$\mathfrak{G}\big(\tau^{\leq\nu}_{c}(\mathrm{ind}_{c}{\mathbb{V}}^{b}_{c}({\mathcal{P}}))\big)$ $=\tau^{\leq-\rho+l\nu}_{r}\circ{\mathrm{I}}^{\mathrm{hb}}_{1}\circ\mathrm{ind}_{c}({\mathbb{V}}^{b}_{c}({\mathcal{P}}))$
$=\tau^{\leq-\rho+l\nu}_{r}\circ\mathrm{ind}_{r}\big(\mathfrak{G}^{b}({\mathbb{V}}^{b}_{c}({\mathcal{P}}))\big)$
$=\tau^{\leq-\rho+l\nu}_{r}(\mathrm{ind}_{r}{\mathbb{V}}^{b}_{r}(P)).$
It follows that $\mathfrak{G}\big({\mathbb{V}}_{c}(\tau^{\leq^{\frac{\infty}{2}}t(\nu)}_{c}{\mathcal{Q}})\big)={\mathbb{V}}_{r}(\tau^{\leq^{\frac{\infty}{2}}t(\nu)}_{r}Q)$. By Lemmas 5.9(1) and 5.18(1), we obtain identifications
${\mathfrak{G}}\circ{\mathbb{V}}_{c}\big(\mathrm{Proj}({\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)})\big)={\mathbb{V}}_{r}\big(\mathrm{Proj}({\mathcal{O}}^{0,\leq^{\frac{\infty}{2}}t(\nu)}_{S})\big),\quad\forall\nu\in\Lambda,$
@@
$\mathrm{Proj}({\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}){\ \xrightarrow{\sim}\ }\mathrm{Proj}({\mathcal{O}}^{0,\leq^{\frac{\infty}{2}}t(\nu)}_{S}),\quad\forall\nu\in\Lambda,$
@@
${\mathfrak{F}}:\ {\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}{\ \xrightarrow{\sim}\ }{\mathcal{O}}^{0}_{S}.$
[…]
@@
${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$${\mathcal{O}}^{0}_{S}$${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}.$$\scriptstyle{{\mathfrak{F}}}\scriptstyle{\simeq}$$\scriptstyle{\mathrm{for}}$$\scriptstyle{\mathrm{for}}$$\scriptstyle{{\mathfrak{F}}^{b}}\scriptstyle{\simeq}$
Hence ${\mathfrak{F}}({\mathcal{M}}_{x})$ ($x\in W_{\mathrm{ex}}$) is isomorphic to $M(x\bullet_{l}0)_{S}$ as a $\Lambda$-graded ${\mathfrak{U}}^{b}_{\zeta}\otimes S$-module. It forces that ${\mathfrak{F}}({\mathcal{M}}_{x})=M(x\bullet_{l}0)_{S}$ as $\Lambda$-graded $U^{\mathrm{hb}}_{\zeta}\otimes S$-modules. Now the assertion follows from (4.21). ∎
@@
The assumption that $l$ is a prime power is used in the results in §3.2 concerning the localization theorem for ${\mathfrak{U}}_{\zeta}$. However, it is expected that this assumption is not necessary (see a proof for type $A$ in [40]). For Theorem 5.23, one may remove this assumption by showing that the equivalence class of the category ${\mathcal{O}}^{0}_{S}$ is independent of $\zeta$ (note that similar independence for small quantum groups appeared in [1].)
Let $\pi_{{}_{\widetilde{{\mathcal{N}}}}}:\widetilde{{\mathcal{N}}}=G\times^{B}{\mathfrak{n}}\rightarrow{\mathfrak{g}}$ be the Springer resolution. We set $\mathsf{A}$$^{\prime}$ be the specialization of $\mathsf{A}$} at $0\in{\mathfrak{t}}$, which satisfies an equivalence [11]
$D^{\mathrm{b}}\mathrm{Coh}(\widetilde{{\mathcal{N}}})\ \simeq\ D^{\mathrm{b}}({\mathsf{A}}^{\prime}\text{-}\mathrm{mod}),$
and therefore is called the non-commutative Springer resolution. Recall that ${\mathsf{A}}_{0}$ is the specialization of $\mathsf{A}$} at $0\in{\mathfrak{g}}$. We abbreviate ${\mathcal{O}}^{\chi_{0}}_{\mathbb{C}}$ as the full subcategory of modules in $\mathcal{O}$$_{\mathbb{C}}$ on which the action of $Z_{\mathrm{HC}}$ factors through $\chi_{0}:Z_{\mathrm{HC}}\rightarrow{\mathbb{C}}$.
@@
(5.12) $\mathrm{Coh}^{G}(\pi_{{}_{\widetilde{{\mathcal{N}}}}}^{*}{\mathsf{A}})={\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}{\ \xrightarrow{\sim}\ }{\mathcal{O}}_{\mathbb{C}}^{0},$
@@
$\mathrm{Coh}^{G}(\pi_{{}_{\widetilde{{\mathcal{N}}}}}^{*}{\mathsf{A}}^{\prime}){\ \xrightarrow{\sim}\ }{\mathcal{O}}_{\mathbb{C}}^{{\chi_{0}}},\quad\text{and}\quad{\mathsf{A}}_{0}\text{-}\mathrm{mod}^{B}{\ \xrightarrow{\sim}\ }{\mathcal{O}}^{u_{\zeta}U^{+}_{\zeta},0}_{\mathbb{C}}.$
Remark 5.26.
Our construction can be applied to obtain a new approach to a similar equivalence for modular analogue of BGG category $\mathcal{O}$}, which is previously established by Losev [25, Equ (1.3) in Thm 1.7]. To be more precise, assume moreover that $l$ is prime and set $\mathbb{F}$} be an algebraically closed field of characteristic $l$. Let $G_{\mathbb{F}}$ be a connected semisimple algebraic group over $\mathbb{F}$}, with a Borel subgroup $B_{\mathbb{F}}$ and a Cartan subgroup $T_{\mathbb{F}}$. Let $N_{\mathbb{F}}$ be the unipotent radical of $B_{\mathbb{F}}$, and let ${\mathfrak{g}}_{\mathbb{F}}=\mathrm{Lie}(G_{\mathbb{F}})={\mathfrak{n}}^{-}_{\mathbb{F}}\oplus{\mathfrak{t}}_{\mathbb{F}}\oplus{\mathfrak{n}}_{\mathbb{F}}$ be the corresponding triangular decomposition with ${\mathfrak{t}}_{\mathbb{F}}=\mathrm{Lie}(T_{\mathbb{F}})$ and ${\mathfrak{n}}_{\mathbb{F}}=\mathrm{Lie}(N_{\mathbb{F}})$. The modular analogue of BGG category $\mathcal{O}$} is the category
${\mathcal{O}}_{\mathbb{F}}:={\mathfrak{g}}_{\mathbb{F}}\text{-}\mathrm{mod}^{B_{\mathbb{F}}}$
of $B_{\mathbb{F}}$-equivariant finitely generated ${\mathfrak{g}}_{\mathbb{F}}$-modules on which the actions of ${\mathfrak{b}}_{\mathbb{F}}$ by restriction of ${\mathfrak{g}}_{\mathbb{F}}$-action and by differential of $B_{\mathbb{F}}$-action coincide. The category ${\mathcal{O}}_{\mathbb{F}}$ can be interpreted as the category $\mathcal{O}$} (with integral weights) associated to the modular analogue $U^{\mathrm{hb}}_{\mathbb{F}}$ of $U^{\mathrm{hb}}_{\zeta}$, which is an algebra with a triangular decomposition
$U^{\mathrm{hb}}_{\mathbb{F}}=U{\mathfrak{n}}^{-}_{\mathbb{F}}\otimes U{\mathfrak{t}}_{\mathbb{F}}\otimes\mathrm{Dist}(N_{\mathbb{F}}),$
where $\mathrm{Dist}(N_{\mathbb{F}})$ is the distribution algebra of $N_{\mathbb{F}}$. Let ${\mathsf{A}}_{\mathbb{F}}$ be the non-commutative Grothendieck resolution of ${\mathfrak{g}}^{(1)}_{\mathbb{F}}$ (the Frobenius twist of ${\mathfrak{g}}_{\mathbb{F}}$). Let $\pi_{\widetilde{{\mathcal{N}}}^{(1)}_{\mathbb{F}}}:\widetilde{{\mathcal{N}}}^{(1)}_{\mathbb{F}}\rightarrow{\mathfrak{g}}^{(1)}_{\mathbb{F}}$ be the Springer resolution. Then there is an equivalence of categories
$\mathrm{Coh}^{G_{\mathbb{F}}^{(1)}}(\pi^{*}_{\widetilde{{\mathcal{N}}}^{(1)}_{\mathbb{F}}}{\mathsf{A}}_{\mathbb{F}})\ \simeq\ {\mathcal{O}}^{0}_{\mathbb{F}},$
where ${\mathcal{O}}^{0}_{\mathbb{F}}$ is the principal block of ${\mathcal{O}}_{\mathbb{F}}$. To adapt our construction to obtain the equivalence above, one may replace Tanisaki’s localization theorem in §3.2 by the result for $U{\mathfrak{g}}_{\mathbb{F}}$ by Bezrukavnikov–Mirković–Rumynin [10] and replace Bezrukavnikov’s equivalence (used in §5.2.4) by its modular analogue recently proved by Bezrukavnikov–Riche [13].